Carson Trucking is considering whether to expand its regional service center in​ Mohab, UT. The expansion requires the expenditure of $11,000,000 on new service equipment and would generate annual net cash inflows from reduced costs of operations equal to ​$4,000,000 per year for each of the next 9 years. In year 9 the firm will also get back a cash flow equal to the salvage value of the​ equipment, which is valued at ​$1 million. ​ Thus, in year 9 the investment cash inflow totals $5,000,000. Calculate the​ project's NPV using a discount rate of10 percent.
Answer #1

NPV=-Initial investment+Annual cash flow/rate*(1-1/(1+rate)^t)+After tax salvage value/(1+rate)^t=-11000000+4000000/10%*(1-1/1.1^9)+1000000/1.1^9=12460192.88